### v1

# On Landis conjecture for positive Schrödinger operators on graphs

Abstract.
In this note we study Landis conjecture for positive Schrödinger operators on graphs. More precisely, we give a decay criterion that ensures when $\mathcal{H}$-harmonic functions for a positive Schrödinger operator $\mathcal{H}$ with potentials bounded from above by $1$ are trivial. We then specifically look at the special cases of $\mathbb{Z}^{d}$ and regular trees for which we get explicit decay criterion. Moreover, we consider the fractional analogue of Landis conjecture on $\mathbb{Z}^{d}$. Our approach relies on the discrete version of Liouville comparison principle which is also proved in this article.
2010 Mathematics Subject Classification. Primary 35J10; Secondary 35B53, 35R02, 39A12.
Keywords: Agmon ground state, Landis Conjecture, Liouville comparison principle, unique continuation at infinity.
1. Introduction
In the late 1960’s Landis conjectured that a $\mathcal{H}$-harmonic function $u$ of a Schrödinger operator $\mathcal{H}=\Delta+V$ on $\mathbb{R}^{d}$, i.e.,
$\mathcal{H}u=\Delta u+Vu:=-\sum_{j=1}^{d}\partial_{j}^{2}u+Vu=0\ \ \ \mbox{on}\ \mathbb{R}^{d}$
with $|V|\leq 1$ which satisfies
$|u(x)|\leq\exp({-C|x|})$
for sufficiently large $C>0$ is trivially zero, [L71, KL88]. Landis also conjectured a weaker version which states that if
$|u(x)|\leq\exp({-|x|^{1+\varepsilon}})$
for $\varepsilon>0$, then $u=0$ in ${\mathbb{R}}^{d}$. Such a result can be seen as a property of unique continuation at infinity for $\mathcal{H}$-harmonic functions in ${\mathbb{R}}^{d}$.
In 1992, Meshkov [M92] provided a counterexample for a complex valued potential in two dimension by constructing a complex valued bounded potential $V$ and a nontrivial $\mathcal{H}$-harmonic function $u$ in ${\mathbb{R}}^{2}$ which satisfies $|u(x)|\leq\exp({-C|x|^{4/3}})$ for some constant $C>0$. Furthermore, Meshkov showed that if
$\sup_{x\in\mathbb{R}^{d}}|u(x)|\exp({C|x|^{4/3}})<\infty$
for all $C>0$, then $u=0$ in ${\mathbb{R}}^{d}$. The result with exponent $2$ instead of $4/3$ already appears in the earlier work [FHHH82] (put $\beta=1$ in Theorem 2.3) which however seems to be neglected in the literature.
Although, due to Meshkov’s results, the Landis conjecture is settled for complex-valued bounded potentials, it still remains open in the real-valued case. There is plenty of research in $\mathbb{R}^{2}$ [DKW17, KSW15, LMNN20]; our bibliography refers only to a small fraction of it. A few results addressed the conjecture in $\mathbb{R}^{d}$ [ABG19, BK05, SS21] and the very recent paper [DP24]. See also in the case of an exterior domain in ${\mathbb{R}}^{d}$ [R21, DP24], a half cylinder in $\mathbb{R}^{d}$ [FK23]. To the best of our knowledge, the strongest results so far available is only for $\mathbb{R}^{2}$ in [LMNN20]. By strongest we mean that the authors assumed the weakest (with compare to the available literature) decay condition on $u$ for the validity of the Landis conjecture. Nevertheless, in ${\mathbb{R}}^{d}$, with additional positivity assumption on $\mathcal{H}$, a sharp decay criterion is recently obtained in [DP24]. One may note that some of the mentioned articles concern more general operators than $\mathcal{H}$. In the Euclidean settings, Landis-type results have been also studied for Dirac operators [C22], the time-dependent Schrödinger equation and the heat equations [EKPV10, EKPV16], see also the references therein. Further, there are results for time-independent and time dependent Schrödinger equations on graphs with discrete Laplacian in [BMP24, BP24, FB19, FV17, JLMP18, LM18]. In $\mathbb{Z}^{d}$, for a bounded potential $V$, it has been shown that if
${\liminf_{N\rightarrow\infty}\frac{1}{N}\log\max_{|x|_{\infty}\in[N,N+1]}|u(x)|<-\|V\|_{\infty}-4d+1}\,,$
then $u=0$ in $\mathbb{Z}^{d}$ [LM18]. We also refer to [FBRS24], where the authors recently studied the Landis-type uniqueness results in a mesh $(h\mathbb{Z})^{d}$, $h>0$ under certain summation criteria on $u$, and analyzed the behaviour of the solutions as the mesh-size $h$ decreases to zero.
The purpose of this paper is to study Landis conjecture on general discrete graphs under the additional assumption that the Schrödinger operator in question is positive in the sense of the quadratic form. For the precise definition of $\mathcal{H}$ on discrete graphs and for more details, see Section 2. Our approach is inspired by the recent paper [DP24], where the corresponding problem in the continuum is studied. On the one hand, the positivity of the operator is clearly a restriction. However, most of the known results in ${\mathbb{R}}^{d}$ effectively make also an assumption of this kind (see for example [ABG19, KSW15, SS21]).
In Theorem 3.1, we show that the $\mathcal{H}$-harmonic functions, i.e., solutions of the equation $\mathcal{H}u\!=\!0$ on a discrete graph, are trivial under much weaker decay conditions than those discussed above. Moreover, we only assume $V\leq 1$ instead of $|V|\leq 1$.
Next, we give an overview of the decay conditions given in Theorem 3.1 and the overall strategy. For the details and precise definitions, we refer the reader to the next sections. We consider $\mathcal{H}$-harmonic functions $u$ which should satisfy an “a priori estimate” on the whole space which is typically not in $\ell^{2}$ and even might only be in $\ell^{\infty}$. Secondly, a much stronger “decay estimate” must be satisfied but only in one direction of the space. Such a $\mathcal{H}$-harmonic $u$ is then shown to be trivial.
The “a priori estimate” will guarantee that $u$ has a constant sign, and therefore, we may assume that $u$ is positive. If the function we compare $u$ to, was in $\ell^{2}$, then positivity of $u$ is trivial because the $\mathcal{H}$-harmonic function $u$ would then be either trivial or an eigenfunction at the bottom of the spectrum which is known to be strictly positive. To go beyond $\ell^{2}$, we prove and employ a discrete version of a Liouville comparison principle, Theorem 2.1, which goes back to [P07] in the continuum. Specifically, the comparison of a $\mathcal{H}$-subharmonic function $u_{+}$ will be made with a “slowly decaying” Agmon ground state of a related critical operator.
For the “decay estimate”, we compare $u$ with $G_{1}$ which is given as the resolvent of $\mathcal{H}_{1}=\Delta+1$ applied to the delta function at a fixed vertex. This $G_{1}$ is a positive solution of the equation $\mathcal{H}_{1}\varphi=0$ of minimal growth at infinity. If we now assume that the potential satisfies $V\leq 1$, then $u$ is also a positive supersolution of $\mathcal{H}_{1}\varphi=0$, and hence, it cannot decay faster than $G_{1}$ in any direction unless $u=0$. For general graphs, this is comprised in Theorem 3.1 and Corollaries 3.33.4 and 3.5. Moreover, Theorem 3.1 is sharp, see the remark below the proof of Theorem 3.1.
While for general graphs, we have of course no explicit estimates of the resolvent, on $\mathbb{Z}^{d}$ and regular trees one has a rather good understanding of $G_{1}$, [MS22, MY12]. Specifically, for a $\mathcal{H}$-harmonic function $u$ of a positive Schrödinger operator $\mathcal{H}$ on $\mathbb{Z}^{d}$ with potential $V\leq 1$, we can show the following: For $d=1,2$ the “a priori estimate" make us impose the condition that $u$ is bounded in the whole space, and for $d\geq 3$ that $u$ has to be bounded by the square root of the Laplacian’s minimal positive Green function $G_{0}$ at 00, which implies that $u\in O(|x|^{(1-d)/2})$ near infinity, see e.g. [Woe00]. More precisely, this uses that $G_{0}^{1/2}$ is an Agmon ground state for $\Delta-W$ for a specific critical Hardy weight $W$, [KPP18]. Furthermore, it is known that one can estimate $G_{1}(x)$ by a multiple of $|x|^{(1-d)/2}\mathrm{e}^{-|x|}$, cf. [MY12, MS22]. Hence, on $\mathbb{Z}^{d}$, if $u$ satisfies the a priori estimate $u\in O(|x|^{(1-d)/2})$ and
$\liminf_{x\to\infty}|u(x)||x|^{(d-1)/2}\mathrm{e}^{|x|}=0,$
then $u=0$ by Theorem 4.1.
On $d$-regular trees, we obtain that if $u\in O(|x|^{1/2}d^{-|x|/2})$ and
$\liminf_{x\to\infty}|u(x)|d^{|x|}=0,$
then $u=0$ by Theorem 4.3. Here we use the considerations for optimal Hardy weights on regular trees by [BSV21].
We furthermore consider the Landis-type results for Schrödinger operator involving the fractional Laplacian $\Delta^{\sigma}$ on $\mathbb{Z}^{d}$, $\sigma\in(0,1)$. We use recent estimates on the Green’s function derived in [DER24] to show that if $u\in O\left({|x|^{2\sigma-d}}\right)$ and
$\liminf_{x\to\infty}|u(x)||x|^{2\sigma+d}=0,$
then $u\!=\!0$ $\mathbb{Z}^{d}$ by Theorem 5.1. Moreover, for $d=1$, the a priori bound which is needed can be improved to $O\left(|x|^{(2\sigma-d)/2}\right)$ in Theorem 5.2 due to recent results by [KN23].
The paper is structured as follows. In the next section, we introduce the set-up and prove a Liouville comparison theorem, Theorem 2.1. In Section 3, we study the problem on general graphs and prove the main abstract result, Theorem 3.1 together with its corollaries, Corollaries 3.33.4 and 3.5. In Section 4, we apply these results to $\mathbb{Z}^{d}$ in Theorem 4.1 and 4.2, and to regular trees in Theorem 4.3. Finally, in Section 5 we study the fractional analogue.
2. Set up and a Liouville comparison theorem
In this section, we recall some basic notions and results in criticality theory that are essential for the development of this article. In addition, we prove the discrete analogue of Liouville comparison principle for Schrödinger operators on graphs.
Throughout the paper, we use the following notation and conventions:
  • For a set $A$, we write #$\#A$ to denote its cardinality.
  • For two subsets $A,B$ in a discrete topological space $X$, we write $A\Subset B$ if $A$ is a compact, i.e., finite subset of $B$.
  • If a function $f\geq 0$ is not trivial, then we say $f$ is positive. Moreover, if $f>0$, then we say $f$ is strictly positive.
  • For positive functions $f,g:A\to{\mathbb{R}}$, we write $f\asymp g$ if there is a constant $C>0$ such that $C^{-1}f\leq g\leq Cf$.
  • For two functions $f,g:A\to{\mathbb{R}}$, we denote $f\vee g$ the pointwise maximum of $f$ and $g$.
  • For two functions $f,g:X\to{\mathbb{R}}$ such that $g$ does not vanish outside of a compact set, we denote
    $f\in O(g)\qquad\Longleftrightarrow\qquad\limsup_{x\to\infty}\frac{|f(x)|}{|g(x)|}<\infty,$
    where the limit $x\to\infty$ is taken in the topology of the one point compactification $X\cup\{\infty\}$ of $X$.
Let $X$ be an infinite countable set equipped with the discrete topology. We denote by $C(X)$ the real valued functions on $X$ and by $C_{c}(X)$ the subspace of functions of compact (i.e. finite) support. For a function $f:X\to{\mathbb{R}}$, we write
$\sum_{X}f=\sum_{x\in X}f(x)$
whenever the sum is absolutely convergent. Any strictly positive function $m:X\to(0,\infty)$ extends to a measure of full support on $X$ via $m(A)=\sum_{x\in A}m(x)$.
A graph $b$ over the measure space $(X,m)$ is a symmetric function $b:X\times X\to[0,\infty)$ with zero diagonal which is locally summable, i.e.,
$\sum_{y\in X}b(x,y)<\infty,\qquad{x\in X}.$
Given a graph $b$, we denote $x\sim y$ whenever $b(x,y)>0$ and think of $x$ and $y$ to be connected by an edge. The (vertex) degree of $x\in X$ is the number of vertices connected to $x$.
Throughout this paper, we assume that the graph $b$ is connected, i.e., for every $x,y\in X$ there are vertices $x=x_{0}\sim\ldots\sim x_{n}=y$ connecting $x$ and $y$ by a path.
We introduce the (positive) Laplacian $\Delta=\Delta_{b,m}$ acting on its formal domain $\mathcal{F}=\mathcal{F}_{b}$
=$\mathcal{F}=\{f:X\to\mathbb{R}\mid\sum_{y\in X}b(x,y)|f(y)|<\infty\mbox{ for all }x\in X\}$
as
$\Delta f(x)=\frac{1}{m(x)}\sum_{y\in X}b(x,y)(f(x)-f(y)).$
For a potential $V:X\to{\mathbb{R}}$, we denote the corresponding Schrödinger operator $\mathcal{H}=\mathcal{H}_{b,V,m}$ acting on $f\in\mathcal{F}$ as
$\mathcal{H}f(x)=\Delta f(x)+\frac{V(x)}{m(x)}f(x).$
A function $u$ is $\mathcal{H}$-$(\overset{super}{sub})harmonic$ on $X$ if $u\in\mathcal{F}$ and $\mathcal{H}u=0$ ($\mathcal{H}u\overset{\geq}{\leq}0$) on $X$. We write $\mathcal{H}\geq 0$ in $X$ if there exists a positive $\mathcal{H}$-superharmonic function $u$ on $X$, and in this case we say that the corresponding Schrödinger operator $\mathcal{H}$ is positive on $X$.
Remark.
We recall that in the continuum the $\mathcal{H}$-(super or sub)harmonic functions in ${\mathbb{R}}^{d}$ are defined in the weak sense i.e., $u\in W^{1,2}_{\mathrm{loc}}({\mathbb{R}}^{d})$ is called an $\mathcal{H}$-(super or sub)harmonic function if
$\langle\mathcal{H}u,\varphi\rangle=\int_{{\mathbb{R}}^{d}}\nabla u\nabla\varphi\ {\mathrm{d}}x+\int_{{\mathbb{R}}^{d}}Vu\varphi\ {\mathrm{d}}x\overset{(\geq\mbox{ or }\leq)}{=}0$
for all $\varphi\in C_{c}^{\infty}({\mathbb{R}}^{d})$ (with $\varphi\geq 0$.) However, in the discrete setting weakly $\mathcal{H}$-harmonic functions are already $\mathcal{H}$-harmonic functions, see [HKLW12, Theorem 2.2 and Corollary 2.3]. Note that, for consistency with the discrete settings, we consider the (positive) Laplacian on ${\mathbb{R}}^{d}$ as $\Delta_{{\mathbb{R}}^{d}}:=-\sum_{j=1}^{d}\partial_{j}^{2}$.
By Green’s formula [KPP20, Lemma 2.1], $\mathcal{H}$ is related to its corresponding energy form $\mathcal{Q}=\mathcal{Q}_{b,V,m}$
$\mathcal{Q}({\varphi})=\frac{1}{2}\sum_{X\times X}b|\nabla{\varphi}|^{2}+\sum_{X}mV{\varphi}^{2},$
which takes finite values for ${\varphi}\in C_{c}(X)$ and where $\nabla_{xy}{\varphi}={\varphi}(x)-{\varphi}(y)$. Unless stated otherwise, we assume throughout the paper that $\mathcal{Q}$ is positive on $C_{c}(X)$, i.e.,
$\mathcal{Q}({\varphi})\geq 0\qquad\mbox{for all }{\varphi}\in C_{c}(X).$
By the Agmon-Allegretto-Piepenbrink-type theorem [KPP20, Theorem 4.2], $\mathcal{Q}({\varphi})\!\geq\!0$ on $C_{c}(X)$ if and only if $\mathcal{H}\geq 0$ in $X$.
A positive function $\phi\!\in\!\mathcal{F}$ is called a positive solution of minimal growth at infinity in $X$ for $\mathcal{H}$, if $\phi>0$ and $\mathcal{H}\phi=0$ in $X\setminus K_{0}$ for some compact $K_{0}\Subset X$, and for any positive $\psi\in\mathcal{F}$ such that $\psi>0$ and ${\mathcal{H}\psi\geq 0}$ on $X\setminus K$ for $K_{0}\Subset K\Subset X$ which satisfies $\phi\leq\psi$ on $K$, one has $\phi\leq\psi$ in $X\setminus K$.
A positive $\mathcal{H}$-harmonic function $\phi$ in $X$ which is a solution of minimal growth at infinity in $X$ is called an Agmon ground state of $\mathcal{H}$.
We call a positive Schrödinger operator $\mathcal{H}$ subcritical if there is a $W\gneq 0$ such that for all ${\varphi}\in C_{c}(X)$
$\mathcal{Q}({\varphi})\geq\sum_{X}mW{\varphi}^{2}.$
We call such a function $W$ a Hardy weight for $\mathcal{H}$. If $\mathcal{H}\geq 0$ is not subcritical, then we say it is critical. Furthermore, we say that a Hardy weight $W$ is a critical Hardy weight for $\mathcal{H}$ if $\mathcal{H}-W$ is critical. If the graph $b$ is connected and the potential $V$ vanishes, i.e., $V=0$, then one also says the graph is transient if $\mathcal{H}$ is subcritical and recurrent if $\mathcal{H}$ is critical.
Recall that the operator $\mathcal{H}$ is critical if and only if there is $x\in X$ and functions ${\varphi}_{n}\in C_{c}(X)$ with ${\varphi}_{n}\!\geq\!0$, $C^{-1}\!\leq\!{\varphi}_{n}(x)\!\leq\!C$ for some $C>0$ and all $n\geq 1$ such that
$\mathcal{Q}({\varphi}_{n})\to 0\quad\mbox{as }n\to\infty,$
cf. [KPP20, Theorem 5.3]. We call such a sequence $({\varphi}_{n})$ a null-sequence for $\mathcal{H}$ with respect to the vertex $x\in X$. Observe that such a null-sequence converges to an Agmon ground state of $\mathcal{H}$, [KPP20, Theorem 5.3]. A further characterization of criticality of $\mathcal{H}$ is that up to scalar multiples there is a unique positive $\mathcal{H}$-superharmonic function which is then $\mathcal{H}$-harmonic and an Agmon ground state, see e.g. [KPP20, Theorem 5.3].
Finally, we recall the ground state transform, cf. [KPP20, Proposition 4.8]. Let $f\in\mathcal{F}$ be a positive function, then for all ${\varphi}\in C_{c}(X)$
$\mathcal{Q}(f{\varphi})=\frac{1}{2}\sum_{X\times X}b(f\otimes f)|\nabla{\varphi}|^{2}+\sum_{X}(f\mathcal{H}f){\varphi}^{2}\,,$
where $f\otimes f:X\times X\to[0,\infty)$ is defined as $f\otimes f(x,y)=f(x)f(y)$.
We now prove a Liouville comparison theorem which is vital for the considerations of this paper, cf. [P07].
Theorem 2.1 (Liouville comparison principle).
Suppose $\mathcal{H}$ and $\mathcal{H}^{\prime}$ are Schrödinger operators associated to connected graphs $b$, $b^{\prime}$ and potentials $V,V^{\prime}$ over $(X,m)$ with positive energy functionals $\mathcal{Q}$ and $\mathcal{Q}^{\prime}$ on $C_{c}(X)$. Let $u\in\mathcal{F}_{b}$ and $v\in\mathcal{F}_{b^{\prime}}$ be such that
  • (a)
    $\mathcal{H}^{\prime}$ is critical and $v>0$ is its Agmon ground state,
  • (b)
    $u_{+}\neq 0$ and $\mathcal{H}u_{+}\leq 0$,
  • (c)
    $b(u_{+}\otimes u_{+})\leq Cb^{\prime}(v\otimes v)$ for some $C>0$.
Then, $\mathcal{H}$ is critical and $u>0$ is an Agmon ground state of $\mathcal{H}$.
Proof.
By (a) there exists a null-sequence $({\varphi}_{k})\subset C_{c}(X)$ with respect to a vertex $x\in X$ for $\mathcal{Q}^{\prime}$. For $k\in\mathbb{N}$, set
$\psi_{k}=\frac{u_{+}{\varphi}_{k}}{v}\,.$
Observe that the ground state transform of $\mathcal{Q}$ with respect to $u_{+}$, (b) and (c) yields
$\mathcal{Q}(\psi_{k})$ $=\frac{1}{2}\sum_{X\times X}b(u_{+}\otimes u_{+})|\nabla({\varphi}_{k}/v)|^{2}+\sum_{X}(u_{+}\mathcal{H}u_{+})({{\varphi}_{k}}/{v})^{2}$
$\leq\frac{C}{2}\sum_{X\times X}b^{\prime}(v\otimes v)|\nabla({\varphi}_{k}/v)|^{2}$
$=C\mathcal{Q}^{\prime}({\varphi}_{k})\to 0,\qquad k\to\infty.$
Thus, $(\psi_{k})$ is a null-sequence for $\mathcal{H}$, and $\mathcal{H}$ is critical. In particular, $(\psi_{k})$ converges pointwise to an Agmon ground state for $\mathcal{H}$. Since $\mathcal{H}^{\prime}$ is critical and $v>0$ is $\mathcal{H}$-harmonic, the null-sequence $({\varphi}_{k})$ converges pointwise to a positive multiple of $v>0$. Hence, $(\psi_{k})$ converges pointwise to a positive multiple of $u_{+}$ which is therefore strictly positive. Thus, $u_{+}=u$ is an Agmon ground state for $\mathcal{H}$. ∎
3. General Landis type theorem
In the present section we prove the main abstract Landis-type theorem and derive some of its important corollaries which will play crucial role in the subsequent section concerning the cases of the Euclidean lattice and regular trees.
Let $\mathcal{H}$ be a Schrödinger operator associated to a connected graph $b$ over $(X,m)$ with potential $V$ and positive energy functional $\mathcal{Q}$.
Since the energy functional of $\Delta$ is positive, the operator $(\Delta+\alpha)$ is subcritical for $\alpha>0$. Hence, for every $o\in X$ there exists a Green function $G_{\alpha}$ which is the smallest strictly positive function $\phi$ such that $(\Delta+\alpha)\phi\geq 1_{o}$, cf. [KPP20, Theorem 5.16]. Moreover, the Green function satisfies
$(\Delta+\alpha)G_{\alpha}=1_{o}$
and is a positive solution of minimal growth at infinity for $(\Delta+\alpha)$ [F24, Theorem 2.5]. Let $L$ be the positive selfadjoint operator associated to the form closure of $\mathcal{Q}_{0}|_{C_{c}(X)}$ in $\ell^{2}(X,m)$ where $\mathcal{Q}_{0}$ is the energy functional of $\Delta$. Then, also by [KPP20, Theorem 5.16]
$G_{\alpha}=(L+\alpha)^{-1}1_{o}\,.$
Moreover, the following limit
[…]
exists but may take the value +$\infty$+ ∞. If $\Delta$ is subcritical, then $G_{0}$ takes only finite (strictly positive) values on $X$ since we assumed the graph to be connected [KLW21].
The next theorem is the abstract main result of the paper.
Theorem 3.1.
Let $\mathcal{H}=\Delta+V$ be a Schrödinger operator associated to a connected graph $b$ over $(X,m)$ with potential $V\leq 1$ and positive energy functional $\mathcal{Q}$. Let $\mathcal{H}^{\prime}$ be a critical Schrödinger operator with Agmon ground state $v>0$ associated to a connected graph $b^{\prime}$ over $(X,m)$. Any $\mathcal{H}$-harmonic function $u$ satisfying
$b(|u|\otimes|u|)\leq Cb^{\prime}(v\otimes v)\ \mbox{ for some }C>0\quad\mbox{and}\quad\liminf_{x\to\infty}\frac{|u(x)|}{G_{1}(x)}=0$
is trivially zero.
For the proof, we need the following lemma which is found for $V\geq 0$ in [KLW21, Lemma 1.9]. Although the proof carries over verbatim to the case of general Schrödinger operators, we include it here for the convenience of the reader.
Lemma 3.2.
The pointwise maximum of two $\mathcal{H}$-subharmonic functions is $\mathcal{H}$-subharmonic.
Proof.
Let $u,v$ be $\mathcal{H}$-subharmonic and $w=u\vee v$. Then, for $x\in X$, we assume $w(x)=u(x)\geq v(x)$ and we have by case distinction for all $y\in X$
$\nabla_{xy}w\leq\nabla_{xy}u.$
This in combination with the assumption $w(x)=u(x)$ for this particular $x$ yields
$\mathcal{H}w(x)=\Delta w(x)+\frac{V(x)}{m(x)}w(x)\leq\Delta u(x)+\frac{V(x)}{m(x)}u(x)=\mathcal{H}u(x)\leq 0.$
For $w(x)=v(x)\geq u(x)$, we obtain analogously $\mathcal{H}w(x)\leq\mathcal{H}v(x)\leq 0$. This settles the claim. ∎
Proof of Theorem 3.1.
Let $u$ be a $\mathcal{H}$-harmonic function, and assume without loss of generality $u_{+}\neq 0$ (otherwise consider $-u$). Then, by Lemma 3.2 above, we have
$\mathcal{H}u_{+}\leq 0.$
We collect the following facts:
  • (a)
    $\mathcal{H}^{\prime}$ is critical and $v>0$ is an Agmon ground state.
  • (b)
    $u_{+}\neq 0$ and $\mathcal{H}u_{+}\leq 0$.
  • (c)
    $b(u_{+}\otimes u_{+})\leq Cb^{\prime}(v\otimes v)$ for some $C>0$.
Thus, by the Liouville comparison theorem, Theorem 2.1, we infer that $\mathcal{H}$ is critical and $u=u_{+}>0$ is an Agmon ground state for $\mathcal{H}$.
Now, we also see that since $V\leq 1$
$(\Delta+1)u\geq(\Delta+V)u=\mathcal{H}u=0.$
As $G_{1}$ is a solution of minimal growth at infinity for $\Delta+1$, we obtain that $\liminf_{x\to\infty}{u(x)}/{G_{1}(x)}=0$ is a contradiction. Thus, any such $\mathcal{H}$-harmonic function is trivially zero. ∎
Remark.
Let us note that Theorem 3.1 is sharp in the following sense. For $V=1-[m(0)/G_{1}(0)]1_{0}\leq 1$, the function $u=G_{1}$ solves
$(\Delta+V)u=(\Delta+1)G_{1}-(1_{0}/G_{1}(0))G_{1}(0)=0.$
In view of the decay conditions assumed on $u$ in the Theorem 3.1, we see the sharpness of our result.
Corollary 3.3.
Let $\mathcal{H}=\Delta+V$ be a Schrödinger operator associated to a connected graph $b$ over $(X,m)$ with potential $V\leq 1$ and positive energy functional $\mathcal{Q}$. Let $W\geq 0$ be such that $\Delta-W$ is critical and let $v$ be its Agmon ground state. If $u\in O(v)$ is a $\mathcal{H}$-harmonic function such that
$\liminf_{x\to\infty}\frac{|u(x)|}{G_{1}(x)}=0,$
then $u=0$.
Proof.
The statement follows directly from the theorem above. ∎
Next, we present two corollaries where the Agmon ground state $v$ associated to a critical Hardy weight $W$ in the above corollary can be replaced by certain exponents of Green function of the Laplacian.
For the next corollary, recall that a function $\phi:X\to(0,\infty)$ is said to be proper if for any compact set $I\subset(0,\infty)$ we have that $\phi^{-1}(I)\subseteq X$ is compact, i.e., finite. Furthermore, we say that $\phi$ is of bounded oscillation if
$\sup_{x\sim y}\frac{\phi(x)}{\phi(y)}<\infty.$
In [KPP18, Theorem 1.1] it is shown for $\mathcal{H}=\Delta$ that a strictly positive $\mathcal{H}$-superharmonic function $\phi$ which is proper, of bounded oscillation and $\mathcal{H}$-harmonic outside of a compact set gives rise to a critical Hardy weight
$W=\frac{\mathcal{H}\phi^{1/2}}{\phi^{1/2}}$
for $\mathcal{H}$. Indeed, this Hardy weight $W$ is even null-critical, i.e., the Agmon ground state ${v=\phi^{1/2}}>0$ of the critical operator $\mathcal{H}-W$ is not in $\ell^{2}(X,Wm)$.
Observe that whenever there is a strictly positive proper function of bounded oscillation, then the graph must be locally finite, i.e., #$\#\{y\in X\mid x\sim y\}<\infty$ for all $x\in X$.
The next corollary allows under the properness and bounded oscillation assumptions on $G_{0}$ to take in the above corollary ${v}=G_{0}^{1/2}$.
Corollary 3.4.
Let $\mathcal{H}\!=\!\Delta+V$ be a Schrödinger operator on the connected graph $b$ over $(X,m)$ with potential $V\!\leq\!1$ and positive energy functional $\mathcal{Q}$. Assume further that $\Delta$ is subcritical and $G_{0}$, the minimal positive Green function of $\Delta$, is proper and of bounded oscillation. Let $u\!\in\!O(G_{0}^{1/2})$ be a $\mathcal{H}$-harmonic function satisfying
$\liminf_{x\to\infty}\frac{|u(x)|}{G_{1}(x)}=0.$
Then $u=0$.
Proof.
By the supersolution construction [KPP18, Theorem 1.1], the function
$W=\frac{\Delta G_{0}^{1/2}}{G_{0}^{1/2}}$
is a critical Hardy weight for $\Delta$ and the associated Agmon ground state of $\Delta-W$ is $v=G_{0}^{1/2}$. Hence, the statement follows now directly from Corollary  3.3 above. ∎
Next, we give a corollary, when no explicit critical Hardy weight is available. In this case, we make a stronger decay assumption on $u$ comparing to the corresponding one in the previous corollary.
Corollary 3.5.
Let $\mathcal{H}=\Delta+V$ be a Schrödinger operator associated to a connected graph $b$ over $(X,m)$ with potential $V\leq 1$ and positive energy functional $\mathcal{Q}$. Assume that $u$ is a $\mathcal{H}$-harmonic function satisfying $u\in O(G_{\alpha})$ for some $\alpha\geq 0$ such that $G_{\alpha}<\infty$, and
$\liminf_{x\to\infty}\frac{|u(x)|}{G_{1}(x)}=0.$
Then $u=0$.
Proof.
Clearly, $(\Delta+\alpha)$ for $\alpha>0$ is subcritical since $\Delta$ is positive, and $\Delta$ is subcritical whenever $G_{0}$ takes finite values. Fix such an $\alpha$. Let $o\in X$ and choose $C_{o}>0$ such that $\mathcal{H}^{\prime}:=\Delta+\alpha-C_{o}1_{o}$ is critical (cf. [DKP24, Lemma 4.4]) with an Agmon ground state $v>0$. Clearly, $O(v)=O(G_{\alpha})$ since $G_{\alpha}$ is a solution of minimal growth at infinity for $\mathcal{H}_{\alpha}$. Hence, $u\in O(G_{\alpha})=O(v)$. Thus, the statement follows from Theorem 3.1. ∎
Finally, we formulate a corollary to relate our result with the summability criteria given in [FBRS24].
Corollary 3.6.
Let $\mathcal{H}$ be a Schrödinger operator associated to a connected graph $b$ over $(X,m)$ with potential $V\leq 1$ and positive energy functional $\mathcal{Q}$. Every $\mathcal{H}$-harmonic function $u\in\ell^{2}(X,G_{1}^{-2})$ is trivial.
Proof.
The summability of $G_{1}^{-2}|u|^{2}$ implies $u\in o(G_{1})$, i.e. $\lim_{x\to\infty}|u(x)|/G_{1}(x)=0$. Thus, the statement follows from Corollary 3.5 above. ∎
Remark (Exterior domains).
It is not hard to derive the Landis-type results on exterior domains, i.e., on sets $\tilde{X}=X\setminus K$ for $K\Subset X$. Let $b$ be a graph over $(X,m)$ and $V$ be a given potential. Assume the form ${\mathcal{Q}}$ is positive on $C_{c}(\tilde{X})$. Then, the operator acting as
~$\tilde{\mathcal{H}}=\tilde{\Delta}+D+V_{\tilde{X}}$
where $\tilde{\Delta}=\Delta_{b|_{\tilde{X}\times\tilde{X}},m|_{\tilde{X}}}$, $V_{\tilde{X}}=V|_{\tilde{X}}$ and $D:\tilde{X}\to[0,\infty)$
$D(x)=\frac{1}{m(x)}\sum_{y\in K}b(x,y)$
is positive on $\tilde{X}$. Note that, the restriction of the graph $b$ to $\tilde{X}\!\times\!\tilde{X}$ is not necessarily connected. However, one can deal with every connected component separately. Furthermore, if $K$ is non-empty, then $D$ does not vanish identically on any connected component of $\tilde{X}$ since the graph is connected. Thus, the operator $\tilde{\Delta}+D$ is always subcritical and we denote the Green’s function of $\tilde{\Delta}+D+\alpha$ for $\alpha\geq 0$ by $\tilde{G}_{\alpha}$. Then one can reproduce all results for $\tilde{\mathcal{H}}$-harmonic functions on $\tilde{X}$ by replacing $G_{\alpha}$ with $\tilde{G}_{\alpha}$. Furthermore, note that if the graph is locally finite, then $D$ is finitely supported. In this case, every positive solution of minimal growth at infinity in $\tilde{X}$ is comparable to a positive solution of minimal growth at infinity in $X$.
In a similar fashion, we can replace $\Delta$ on $X$ by an arbitrary operator $\Delta+D$ on $X$ for a positive potential $D:X\to[0,\infty)$ and obtain analogous results.
4. The Euclidean lattice and regular trees
In this section we apply the abstract results of the previous section to graphs where the asymptotics of the Green functions for the relevant Schrödinger operators are known. We first consider the Euclidean lattice and then the case of trees.
4.1. Euclidean lattice
In the continuum setting of $\mathbb{R}^{d}$ the Green function $G_{\mathbb{R}^{d},1}$ of $\Delta_{\mathbb{R}^{d}}+1$, where $\Delta_{\mathbb{R}^{d}}=-\sum_{j=1}^{d}\partial_{j}^{2}$ has the asymptotics [DP24, Appendix]
$G_{\mathbb{R}^{d},1}(x)\asymp|x|^{(1-d)/2}\mathrm{e}^{-|x|}\qquad\mbox{as }x\to\infty,$
For the Euclidean lattice, the situation is substantially more complicated as $\mathbb{Z}^{d}$ lacks the spherical symmetry, and therefore does not allow for the reduction to a one-dimensional problem. However, the asymptotics of $G_{1}$ are still rather well understood by now [MS22, MY12] and we will take advantage of these results.
To be more precise let $X=\mathbb{Z}^{d}$. The Laplacian $\Delta$ with standard weights, i.e., $b(x,y)=1_{|x-y|=1}$ for $x,y\in\mathbb{Z}^{d}$ and $m=1$, acts as
$\Delta f(x)=\sum_{|y-x|=1}(f(x)-f(y))$
on all functions $f:X\to\mathbb{R}$. Furthermore, we denote the Euclidean norm on $\mathbb{Z}^{d}$ by $|x|=(|x_{1}|^{2}+\ldots+|x_{d}|^{2})^{1/2}$. With slight abuse of notation we also write $x$ for the identity function on $\mathbb{Z}^{{}^{d}}$ and $x_{i}$ for the projection on the $i$-th component, $i=1,\ldots,d$.
We prove the following theorem which is a direct consequence of Theorem 3.1 above and the asymptotics obtained in [MS22, MY12].
Theorem 4.1.
Let $u$ be a $\mathcal{H}$-harmonic function of a Schrödinger operator $\mathcal{H}=\Delta+V$ on $\mathbb{Z}^{d}$ with $V\leq 1$ with positive energy functional. If
  • (a)
    $u$ is bounded for $d=1,2$,
  • (b)
    $u$ satisfies $u\in O\left(|x|^{(2-d)/2}\right)$ for $d\geq 3$,
and
$\liminf_{x\to\infty}|u(x)||x|^{(d-1)/2}\mathrm{e}^{|x|}=0,$
then $u=0$.
For the proof we discuss the decay of $G_{\alpha}$ on $\mathbb{Z}^{d}$ which was studied in the literature. In [MS22] it is shown that the function
$C_{a}=\left(\frac{1}{2d}\Delta+a^{2}\right)^{-1}1_{0}\qquad\mbox{on }\ \mathbb{Z}^{d},\ a\in{\mathbb{R}}$
has the asymptotics ${m_{a}^{(d-3)/2}}{|x|_{a}^{-(d-1)/2}}\mathrm{e}^{-m_{a}|x|_{a}}$, where $m_{a}=\cosh^{-1}\left(1+da^{2}\right)$ and
$|x|_{a}:=\frac{1}{m_{a}}\sum_{i=1}^{d}x_{i}\sinh^{-1}\left(x_{i}r(x)\right)$
with $r(x)$ being the unique solution to
$\frac{1}{d}\sum_{i=1}^{d}\sqrt{1+x_{i}^{2}r(x)^{2}}=1+a^{2}.$
Note that a unique solution exists as the left hand side is strictly monotone in $r\geq 0$. In [MY12, Theorem 3.2] it is shown with a somewhat different notation that
$C_{a}(x)=\frac{m_{a}^{(d-3)/2}}{{|x|_{a}^{(d-1)/2}}}\mathrm{e}^{-m_{a}|x|_{a}}(1+o(1)).$
The reason we cite [MY12, Theorem 3.2] for the asymptotics over [MS22, Theorem 1.3] is that the authors of [MS22] consider directional asymptotics $nx$ for $n\to\infty$ for fixed $x$ and variable $a$ rather than estimates which hold for all $x$ but for fixed $a$. However, the benefit of the considerations of [MS22] is that they identify $|\cdot|_{a}$ as a norm. To relate the results of [MY12] and [MS22] to our situation, we observe that for $\alpha>0$
$G_{\alpha}=\left(\Delta+\alpha\right)^{-1}1_{0}=\frac{1}{2d}\left(\frac{1}{2d}\Delta+\frac{\alpha}{2d}\right)^{-1}1_{0}=\frac{1}{2d}C_{a}$
with $a^{2}={\alpha}/{2d}.$
Proof of Theorem 4.1.
For $a^{2}<2$, we have with $a^{2}={\alpha}/({2d})$ that $\alpha<4d$. Thus, we have, confer Lemma A.2 and Lemma A.1,
$m_{a}|x|_{a}\leq\sqrt{2a^{2}d}|x|=\sqrt{\alpha}|x|\leq\sqrt{\alpha}|x|_{a}.$
Thus, we obtain for $a^{2}=1/(2d)$, i.e., $\alpha=1$,
${|x|^{(d-1)/2}}\mathrm{e}^{|x|}\geq C\cdot{|x|_{a}^{(d-1)/2}}\mathrm{e}^{m_{a}|x|_{a}}\geq C\cdot C_{a}(x)^{-1}=C\cdot G_{1}(x)^{-1}.$
Hence, the assumption implies
$\liminf_{x\to\infty}\frac{|u(x)|}{G_{1}(x)}=0.$
For $d=1,2$, to conclude the statement we apply Corollary 3.3. In this case, it is well known that the operator $\Delta$ is critical (i.e, recurrent) and, therefore, $v=1$ is an Agmon ground state. Thus, if $u$ is bounded, then $u\in O(v)$. Hence, $u=0$ by Corollary 3.3.
For $d\geq 3$, one knows that the Green function $G_{0}$ at $\alpha=0$ satisfies $G_{0}\in{O(|x|^{2-d})}$, see e.g. [Woe00, Theorem 25.11], which is proper and of bounded oscillation. Thus, our assumption $u\in O({|x|^{(2-d)/2}})$ ensures that $u\in O(G_{0}^{1/2})$. Hence, Corollary 3.4 implies $u=0$. ∎
We add another variant of the theorem above. This time we focus on the decay on the axis.
Theorem 4.2.
Let $u$ be a $\mathcal{H}$-harmonic function of a Schrödinger operator $\mathcal{H}=\Delta+V$ on $\mathbb{Z}^{d}$ with $V\leq 1$ with positive energy functional, and let $\lambda=\cosh^{-1}(3/2)=0.962...<1$. If
  • (a)
    $u$ is bounded for $d=1,2$,
  • (b)
    $u$ satisfies $u\in O\left({|x|^{(2-d)/2}}\right)$ for $d\geq 3$,
and for an element $e_{j}$, $j=1,\ldots,d$ of the standard basis of $\mathbb{Z}^{d}$
$\liminf_{n\to\infty}|u(ne_{j})|n^{(d-1)/2}\mathrm{e}^{\lambda n}=0,$
then $u=0$.
Proof.
By Lemma A.2, for $x=ne_{j}$ we have that $|x|_{a}=|x_{j}|=|n|$, and $m_{a}=\cosh^{-1}(3/2)$ for $a^{2}=1/(2d)$. Hence, $G_{1}(ne_{j})=C\cdot n^{(1-d)/2}\mathrm{e}^{-\lambda n}(1+o(1))$. Thus, the result follows along the lines of the proof of Theorem 4.1. ∎
Remark.
(a) In view of Corollary 3.6, one can see that the summability criteria on $u$ given in [FBRS24] ensures that the assumptions of Theorem 4.1 are satisfied, and hence $u$ is trivial on $\mathbb{Z}^{d}$.
(b) In [LM18], Lyubarskii and Malinnikova proved that if $u$ is $\mathcal{H}$-harmonic on ${\mathbb{Z}}^{d}$ and satisfies the decay condition
${\liminf_{N\rightarrow\infty}\frac{1}{N}\log\max_{|x|_{\infty}\in\{N,N+1\}}|u(x)|<-\|V\|_{\infty}-4d+1},$
then $u=0$. For $\|V\|_{\infty}\leq 1$, this result also follows from Theorem 4.1.
[…]
$|u(x_{k})||x_{k}|^{\frac{d-1}{2}}\mathrm{e}^{|x_{k}|}\leq|x_{k}|^{\frac{d-1}{2}}\mathrm{e}^{|x_{k}|-(4d-1)|x_{k}|_{\infty}}\leq|x_{k}|^{\frac{d-1}{2}}{\mathrm{e}^{-2|x_{k}|}}\to 0$
as $k\to\infty$. Thus, the liminf condition of Theorem 4.1 is satisfied.
Next, denote $M_{N}=\max_{|x|_{\infty}\in\{N,N-1\}}|u(x)|$ for $N\in{\mathbb{N}}$. We first claim that $M_{N}\leq{\mathrm{e}}^{-N(4d-1+\|V\|_{\infty})}$ for $N$ sufficiently large. Otherwise, there exists $N_{0}\in{\mathbb{N}}$ such that $M_{N_{0}}>{\mathrm{e}}^{-N_{0}(4d-1+\|V\|_{\infty})}$. On the other hand, it has been shown in the proof of [FBRS24, Lemma 7.2] that
$M_{N+1}\geq\frac{M_{N}}{(4d-1+\|V\|_{\infty})}\geq M_{N}{\mathrm{e}}^{-(4d-1+\|V\|_{\infty})}\,.$
It follows that $M_{N_{0}+1}>{\mathrm{e}}^{-(N_{0}+1)(4d-1+\|V\|_{\infty})}$ and by induction,
$M_{N_{0}+n}>{\mathrm{e}}^{-(N_{0}+n)(4d-1+\|V\|_{\infty})},\quad n\in{\mathbb{N}}.$
But this contradicts the above liminf condition of [LM18]. This proves our claim.
Hence, we infer that there exists $C>0$ such that $|u(x)|\leq C{\mathrm{e}}^{-|x|_{\infty}(4d-1+\|V\|_{\infty})}$ on $\mathbb{Z}^{d}$. This in particular gives $|u(x)|\leq C{\mathrm{e}}^{-4\sqrt{d}|x|}$ on $\mathbb{Z}^{d}$ for some $C>0$, and consequently, $u\in O(|x|^{(2-d)/2})$. Thus, by Theorem 4.1 $u=0$.
4.2. Regular trees
We consider a tree with countably infinite vertex set $X$. We fix an arbitrary vertex $o$ and denote by $|x|=d(x,o)$ the combinatorial graph distance to $o$ from a vertex $x$. A tree is said to be regular if the vertex degree is a constant function. Again, we consider the Laplacian $\Delta$ with standard weights $b(x,y)=1_{d(x,y)=1}$ and let $m=1$.
Theorem 4.3.
Every $\mathcal{H}$-harmonic function $u$ of a Schrödinger operator $\mathcal{H}=\Delta+V$ with $V\leq 1$ and positive energy functional on a regular tree with degree equals to $d$ and
$u\in O\left(|x|^{\frac{1}{2}}d^{-\frac{|x|}{2}}\right)\quad\mbox{and}\quad\liminf_{x\to\infty}|u(x)|d^{|x|}=0$
is trivial, i.e. $u=0$.
Proof.
We aim to apply Corollary 3.3. We start by estimating the positive minimal Green function $G_{\alpha}=(\Delta+\alpha)^{-1}1_{o}$ with $\alpha=1$. For a vertex $x$, we consider the forward tree $T_{x}$ of $x$, that is, the subgraph $b_{x}$ of $b$ on the vertex set $X_{x}=\{y\in X\mid d(o,y)=d(o,x)+d(x,y)\}$. Furthermore, we consider the Dirichlet Laplacian $\Delta_{x}^{(D)}$ with respect to the forward tree $T_{x}$ which is $\Delta_{x}^{(D)}=\Delta_{b_{x},1}+1_{x}$. Furthermore, we denote by
$G^{(x)}_{\alpha}=(\Delta_{x}^{(D)}+\alpha)^{-1}1_{x}$
the Green function for the forward tree of $x$. Then, for $x$ and the unique path $o=x_{0}\sim\ldots\sim x_{n}=x$ connecting $x$ and $o$, we have
$G_{\alpha}(x)=G_{\alpha}(o)\prod_{j=1}^{n}G^{(x_{j})}_{\alpha}(x_{j})$
and we also have for a vertex $x$ with number of neighbors $d(x)$
$-\frac{1}{G_{\alpha}^{(x)}(x)}=-\alpha-d(x)+\sum_{y\in X_{x}\setminus\{x\},y\sim x}G_{\alpha}^{(y)}(y)$
see e.g. [ASW06, Equation (2.9) and (2.10)], [Kle98, Proposition 2.1] or [Kel, Proposition 2.7. and Proposition 2.10.]. For a regular tree with degree $d=d(x)$ and $\alpha=1$, we have that all forward Green functions are the same, i.e., $G_{1}^{(x_{j})}(x_{j})=g$ for some $g>0$. Thus, we have to solve
$dg^{2}-({d+1})g+1=0$
which gives
$g_{\pm}=\frac{d+1}{2d}\pm\sqrt{\frac{(d+1)^{2}}{4d^{2}}-\frac{1}{d}}=\frac{d+1\pm(d-1)}{2d}\,.$
Since $G_{1}^{(x_{j})}$ is the smallest positive solution to $(\Delta^{(D)}_{x_{j}}+1)u=1_{x_{j}}$ and $g_{-}=1/d<1=g_{+}$, we conclude $G_{1}^{(x_{j})}(x_{j})=g_{-}={1}/{d}$. Therefore, $G_{1}(x)=G_{1}(o)d^{-|x|}$. Thus,
$\liminf_{x\to\infty}\frac{|u(x)|}{G_{1}(x)}\leq C\liminf_{x\to\infty}{|u(x)|}d^{|x|}=0.$
Secondly, we consider the ground state $v$ associated to a critical Hardy weight for $\Delta$. In the proof of [BSV21, Theorem 2.7], the authors construct a critical Hardy weight $W=\Delta[(\phi_{1}\phi_{2})^{1/2}]/(\phi_{1}\phi_{2})^{1/2}$ for $\Delta$. The functions $\phi_{1}$, $\phi_{2}$ are given for $|x|\geq 1$ by
$\phi_{1}(x)=|x|d^{-|x|/2}\qquad\mbox{and}\qquad\phi_{2}=d^{-|x|/2}$
and $\phi_{1}(o)=\phi_{2}(o)$ are some positive constant. Clearly, the function
$v(x)=(\phi_{1}(x)\phi_{2}(x))^{1/2}=|x|^{1/2}d^{-|x|/2}$
is positive and satisfies
$(\Delta-W)v=\Delta v-\frac{\Delta v}{v}v=0.$
Since $W$ is proven in [BSV21, Theorem 2.7] to be a critical Hardy weight, $v$ is an Agmon ground state of $\Delta-W$. Hence, by assumption $u\in O(v)$. Thus, $u=0$ follows from Corollary 3.3. ∎
5. On the Fractional Laplacian
On the measure space $(X,m)$ consider the Hilbert space $\ell^{2}(X,m)$. For a given graph $b$ over $(X,m)$, let $\mathcal{Q}_{0}$ be the energy functional of $\Delta$, and let $L$ be the positive selfadjoint operator on $\ell^{2}(X,m)$ associated to the form closure of $\mathcal{Q}_{0}|_{C_{c}(X)}$ in $\ell^{2}(X,m)$ (see Section 3).
For $\sigma\in(0,1)$, the discrete fractional Laplacian $L^{\sigma}$ on $\ell^{2}(X,m)$ is defined via the spectral theorem. Then $L^{\sigma}$ satisfies
$L^{\sigma}f(x)=\frac{1}{|\Gamma(-\sigma)|}\int_{0}^{\infty}(I-\mathrm{e}^{-tL})f(x)\frac{\,\mathrm{d}t}{t^{1+\sigma}}\,,$
where $\mathrm{e}^{-tL}$ is the heat semigroup of the Laplacian $L$. Note that the semigroup $\mathrm{e}^{-tL}$ can be extended to $\ell^{\infty}(X)$ and the graph is called stochastically complete if $\mathrm{e}^{-tL}1=1$. In this case it can be observed (see for example [KN23]) that
$L^{\sigma}f(x)$ $=\frac{1}{|\Gamma(-\sigma)|}\int_{0}^{\infty}\left((\mathrm{e}^{-tL}1)f(x)-(\mathrm{e}^{-tL}f)(x)\right)\frac{\,\mathrm{d}t}{t^{1+\sigma}}$
$=\frac{1}{|\Gamma(-\sigma)|}\int_{0}^{\infty}\sum_{y\neq x}(\mathrm{e}^{-tL}1_{y}(x))(f(x)-f(y))\frac{\,\mathrm{d}t}{t^{1+\sigma}}$
$=\frac{1}{|\Gamma(-\sigma)|}\sum_{y\in X}b_{\sigma}(x,y)|(\nabla_{xy}f)$
with
$b_{\sigma}(x,y)=\int_{0}^{\infty}\mathrm{e}^{-tL}1_{y}(x)\frac{\,\mathrm{d}t}{t^{1+\sigma}}$
which again gives rise to a graph Laplacian for a weighted graph $b_{\sigma}$ which we denote by $\Delta^{\sigma}=\Delta_{b_{\sigma},0,m}$. Furthermore, this operator defined on $\mathcal{F}_{b_{\sigma}}$ is an extension of $L^{\sigma}$. Since the semigroup $\mathrm{e}^{-tL}$ maps positive functions to strictly positive functions on connected graphs, $b_{\sigma}$ will be complete, i.e. every two vertices are adjacent. Hence, as the original graph is infinite, the graph $b_{\sigma}$ will be non-locally finite.
In view of the above discussion, we will use the notions $\mathcal{H}^{\sigma}$-harmonic function, Green function $G_{\alpha}^{\sigma}$ of $\Delta^{\sigma}+\alpha$, (sub)criticality of $\Delta^{\sigma}$, critical Hardy weight for $\Delta^{\sigma}$, etc., for the corresponding notions associated with $\Delta_{b_{\sigma},0,m}$. Therefore, the theory developed in Section 3 and in particular the Liouville comparison principle, Theorem 2.1, applies for fractional Schrödinger operators of the form
$\mathcal{H}^{\sigma}=\Delta^{\sigma}+V\,,\ \ \sigma\in(0,1)\,.$
Consequently, we can study the Landis-type unique continuation results for fractional Schrödinger equations.
The global unique continuation property of the fractional Laplacian on a mesh $(h\mathbb{Z})^{d}$ was recently investigated in [FRR24]. It has been noted that although the global unique continuation property holds for the fractional Laplacian in the continuum, it fails in the discrete case. Nevertheless, the authors showed in [FRR24, Proposition 1.2 & 1.4] that certain global unique continuation property holds for $\ell^{2}$-solutions. Moreover, they have also investigated the global unique continuation for $\mathcal{H}^{\sigma}$.
[…]
Theorem 5.1.
Let $\sigma\in(0,1)$ be such that $0<2\sigma<d$, and $u$ be a $\mathcal{H}^{\sigma}$-harmonic function of a fractional Schrödinger operator $\mathcal{H}^{\sigma}=\Delta^{\sigma}+V$ on $\mathbb{Z}^{d}$ with $V\leq 1$ and positive energy functional. If $u$ satisfies
$u\in O\left({|x|^{2\sigma-d}}\right)\quad\mbox{ and }\quad\liminf_{x\to\infty}|u(x)||x|^{2\sigma+d}=0,$
then $u=0$.
Proof.
It is known that the Green function $G_{0}^{\sigma}$ of the fractional Laplacian $\Delta^{\sigma}$ on $\mathbb{Z}^{d}$ behaves as $G_{0}^{\sigma}\asymp|x|^{2\sigma-d}$ when $0<2\sigma<d$, see [DER24, Theorem 8]. Thus, due to our assumption, $u\in O(G_{0}^{\sigma})$.
On the other hand, the Green function $G_{1}^{\sigma}$ of the fractional Laplacian $\Delta^{\sigma}+1$ on $\mathbb{Z}^{d}$ behaves as $G_{1}^{\sigma}\asymp|x|^{-(2\sigma+d)}$ [DER24, Theorem 7].
Therefore, the conclusion $u=0$ follows from Corollary 3.5. ∎
Remark.
[…]
Next we would like to improve the above result by replacing the a priori bound $u\in O\left({|x|^{2\sigma-d}}\right)$ in this theorem above by the weaker bound $u\in O(|x|^{(2\sigma-d)/2})$. We may think of constructing a critical Hardy weight $W_{\sigma}$ for $\Delta^{\sigma}$ and show that the Agmon ground state $v$ of $\Delta^{\sigma}-W_{\sigma}$ satisfies $v\asymp|x|^{(2\sigma-d)/2}$. Having the Green function $G_{0}^{\sigma}$ and its explicit asymptotic, one may anticipate to use similar supersolution construction as in [KPP18, Theorem 1.1.] to construct an optimal Hardy weight $W_{\sigma}$ and get the desired asymptotic of $v$. This is not immediate, since the graph $b_{\sigma}$ is non-locally finite and hence $G_{0}^{\sigma}$ fails to satisfy the bounded-oscillation property, which is required for the optimality result. However, in one dimension, due to the recent results in [KN23], we have an optimal Hardy weight $W_{\sigma}$ of $\Delta^{\sigma}$ on $\mathbb{Z}$ and we know the asymptotic of the associated Agmon ground state. Thus, we have the following result.
Theorem 5.2.
Let $\sigma\in(0,1/2)$ and $d=1$. Assume that $u$ is a $\mathcal{H}^{\sigma}$-harmonic function of a fractional Schrödinger operator $\mathcal{H}^{\sigma}=\Delta^{\sigma}+V$ on $\mathbb{Z}$ with $V\leq 1$ and positive energy functional. If
$u\in O\left({|x|^{(2\sigma-d)/2}}\right)\quad\mbox{ and }\quad\liminf_{x\to\infty}|u(x)||x|^{1+2\sigma}=0,$
then $u=0$.
Proof.
Using [KN23, Theorem 1 & 5], there exists an optimal Hardy weight $W_{\sigma}$ such that the associated Agmon ground state $v$ of $\Delta^{\sigma}-W_{\sigma}$ on $\mathbb{Z}$ satisfies $v\asymp|x|^{(2\sigma-d)/2}$. Thus, due to our assumption, $u\in O(v)$.
On the other hand, we also have that Green function $G_{1}^{\sigma}$ of the fractional Laplacian $\Delta^{\sigma}+1$ on $\mathbb{Z}$ behaves as $G_{1}^{\sigma}\asymp|x|^{-(1+2\sigma)}$ [DER24, Theorem 7]. Hence, Corollary 3.3 implies that $u=0$ on $\mathbb{Z}$. ∎
Appendix A Estimates of norms on $\mathbb{Z}^{d}$ for the resolvent asymptotics
In [MS22] one finds the following estimate of the $|\cdot|_{a}$ norm by the $\ell^{2}$-norm ||$\cdot$|| ⋅ | and the $\ell^{1}$-norm $\|\cdot\|_{1}$ on $\mathbb{Z}^{d}$.
Lemma A.1 (Proposition 1.2. in [MS22]).
For all $a>0$, we have
$|x|\leq|x|_{a}\leq\|x\|_{1}\,.$
Moreover, we also get the following $\ell^{2}$-upper bound for $|x|_{a}$
Lemma A.2.
For $a^{2}<2$, we have
$m_{a}|x|_{a}\leq\sqrt{2a^{2}d}|x|\,.$
Furthermore, for any element $x$ of an axis $\{(0,\ldots,x_{j},\ldots,0)\in\mathbb{Z}^{d}\mid x_{j}\in\mathbb{Z}\}$, $j=1,\ldots,d$, we have
$|x|_{a}=|x_{j}|\,.$
Proof.
We start with the second statement and consider $x=(x_{1},0,\ldots,0)\neq 0$. For any such element on the axis, the equation $\frac{1}{d}\sum_{i=1}^{d}\sqrt{1+x_{i}^{2}r(x)^{2}}=1+a^{2}$ translates to $d-1+\sqrt{1+x_{1}^{2}r(x)^{2}}=d(1+a^{2})$ and, thus,
$r(x)=\frac{1}{|x_{1}|}\sqrt{(1+da^{2})^{2}-1}\,.$
Hence, applying the hyperbolic trigonometric formula, $\cosh^{2}-\sinh^{2}=1$, yields
$|x|_{a}\!$ $=|x_{1}|\frac{\sinh^{-1}\!\left(\!\sqrt{(1+da^{2})^{2}-1}\right)}{\cosh^{-1}(1+da^{2})}\!=\!|x_{1}|\frac{\sinh^{-1}\!\left(\!\sqrt{\cosh^{2}\left(\cosh^{-1}(1+da^{2})\!\right)-1}\right)}{\cosh^{-1}(1+da^{2})}$
$=|x_{1}|\,.$
To obtain the estimate for arbitrary elements, we determine the critical points of the function
$F(x,r)=\sum_{i=1}^{d}x_{i}\sinh^{-1}\left(x_{i}r\right)$
for $x\in{\mathbb{Z}}^{d}$, $r\in[0,\infty)$, under the constraints
$f(x,r)=\sum_{i=1}^{d}\sqrt{1+x_{i}^{2}r^{2}}-(1+a^{2})d=0,\qquad\mbox{and}\qquad g(x)=\frac{1}{2}\left(\sum_{i=1}^{d}x_{i}^{2}-1\right)=0,$
since $m_{a}|x|_{a}=F(x,r(x))$, $f(x,r(x))=0$ is the defining equation for $r(x)$ and $|x|=1$ is equivalent to $g(x)=0$. To apply the methods of Lagrange multipliers, we compute the partial derivatives $\partial_{i}=\partial_{x_{i}}$ and $\partial_{r}$
$\partial_{i}F(x,r)$ $=\sinh^{-1}\left(x_{i}r\right)+\frac{x_{i}r}{\sqrt{1+x_{i}^{2}r^{2}}}\,,\quad$ $\partial_{r}F(x,r)=\sum_{i=1}^{d}\frac{x_{i}^{2}}{\sqrt{1+x_{i}^{2}r^{2}}}\,,$
$\partial_{i}f(x,r)$ $=\frac{x_{i}r^{2}}{\sqrt{1+x_{i}^{2}r^{2}}}\,,\quad$ $\partial_{r}f(x,r)=\sum_{i=1}^{d}\frac{x_{i}^{2}r}{\sqrt{1+x_{i}^{2}r^{2}}}\,,$
$\partial_{i}g(x)$ $=x_{i}\,,\quad$ $\partial_{r}g(x)=0.$
Hence, letting $f=g=0$ and for $\lambda,\mu\in{\mathbb{R}}$, we solve
(I) $\qquad\partial_{i}F+\lambda\partial_{i}f+\mu\partial_{i}g$ $=0,$
(II) $\qquad\partial_{r}F+\lambda\partial_{r}f$ $=0.$
The second equation (II) reads
$\sum_{i=1}^{d}\frac{x_{i}^{2}}{\sqrt{1+x_{i}^{2}r^{2}}}(1+\lambda r)=0$
so either $x_{i}=0$ for all $i$ or
$\lambda=-\frac{1}{r}.$
Now, $g(x)=0$ implies that not all $x_{i}$ can vanish and therefore equation (I) yields
$\sinh^{-1}\left(x_{i}r\right)+\mu x_{i}=0.$
Hence, $x_{i}=0$ for possibly some $i=1,\ldots,d$ and all other $x_{i}$ have to agree (since $\sinh^{-1}$ and the identity are both monotone increasing functions). We infer from this and $g(x)=0$ that
$|x_{i}|=|x_{j}|=\frac{1}{\sqrt{k}}$
for all $i,j\in\{l\in\{1,\ldots,d\}\mid{x_{l}}\neq 0\}=:K$ and $k=\#K$. Thus, we conclude that the critical points are of the form
$\frac{1}{\sqrt{k}}(\sigma_{1}1_{K}(1),\ldots,\sigma_{d}1_{K}(d)),$
with $\sigma_{1},\ldots,\sigma_{d}\in\{\pm 1\}$. Computing $r$ for these critical points from $f(x,r)=0$, we obtain
$(1+a^{2})d=\sum_{i\in K}\sqrt{1+{r^{2}}/{k}}+(d-k)=k\sqrt{1+{r^{2}}/{k}}+(d-k).$
Hence,
$\left(1+{a^{2}d}/{k}\right)^{2}=1+{r^{2}}/{k}.$
Therefore,
$r=\sqrt{\left(\left(1+{a^{2}d}/{k}\right)^{2}-1\right)k}\,.$
Plugging the critical points $(x,r)$ into $F$, we obtain since $\sinh^{-1}$ is an odd function
$F(x,r)$ $=\sum_{i\in K}\frac{1}{\sqrt{k}}\sinh^{-1}\left(\sqrt{\left(\left(1+{a^{2}d}/{k}\right)^{2}-1\right)}\right)$
$={\sqrt{k}}\sinh^{-1}\left(\sqrt{\left(\left(1+{a^{2}d}/{k}\right)^{2}-1\right)}\right)$
and applying the hyperbolic trigonometric formula $\cosh^{2}-\sinh^{2}=1$ yields
$\ldots$ $={\sqrt{k}}\sinh^{-1}\left(\sqrt{\left(\cosh^{2}\left(\cosh^{-1}\left(1+{a^{2}d}/{k}\right)\right)-1\right)}\right)$
$={\sqrt{k}}\sinh^{-1}\left(\sinh\left(\cosh^{-1}\left(1+{a^{2}d}/{k}\right)\right)\right)$
$={\sqrt{k}}\cosh^{-1}\left(1+{a^{2}d}/{k}\right):=G(k).$
By expanding $\cosh^{-1}$ into a power series for $|t|<2$
$\cosh^{-1}(1+t)=\sqrt{2t}\sum_{n=0}^{\infty}\frac{2^{-n}(-t)^{n}{(1/2)^{(n)}}}{{(2n+1)n!}}\,,$
where the Pochhammer symbol ${s^{(n)}}$ is given by $s^{(n)}=s(s+1)\ldots(s+n-1)$, (it includes $n$ factors). We obtain for $t<2$ and $n\in 2\mathbb{N}+1$
$-\frac{2^{-n}t^{n}(\tfrac{1}{2})^{(n)}}{(2n+1)n!}+\frac{2^{-n-1}t^{n+1}(\tfrac{1}{2})^{(n+1)}}{(2n+3)(n+1)!}$ $=-\frac{2^{-n}t^{n}(\tfrac{1}{2})^{(n)}}{(2n+1)n!}\!\left(\!1\!-\!\frac{t(\tfrac{1}{2}+n)(2n+1)n!}{2(2n+3)(n+1)!}\!\right)\!\leq\!0.$
Thus, since $a^{2}d/k\leq a^{2}<2$, we can estimate the series by the first summand and obtain under the constraints $f=g=0$
$F(x,r)\leq G(k)\leq\sqrt{2a^{2}d}\,.$
From this, we infer for $|x|=1$
$m_{a}|x|_{a}=F(x,r(x))\leq\sqrt{2a^{2}d}\,,$
and thus, as $|\cdot|_{a}$ is a norm, $m_{a}|x_{a}|\leq\sqrt{2a^{2}d}|x|$. This finishes the proof. ∎
Acknowledgments: The authors gratefully acknowledge the financial support of the DFG. U.D. and Y.P. acknowledge the support of the Israel Science Foundation (grant 637/19) founded by the Israel Academy of Sciences and Humanities. U.D. is also supported in part by a fellowship from the Lady Davis Foundation. The authors also want to thank Rupert Frank and Dirk Hundertmark for the helpful references to the literature as well as Ira Herbst and Richard Froese for explanations on their earlier results.